OpenStudy (otonogold):

Multiple Choice :) Which graph best represents the feasibility region for the system shown above? x≥1 x≤7 y≥2 y≤ - 1/3x + 6

OpenStudy (otonogold):

A

OpenStudy (otonogold):

B

OpenStudy (otonogold):

C

OpenStudy (otonogold):

D

OpenStudy (anonymous):

I have this same question, it's actually y is greater than or equal to 2, not x!

OpenStudy (otonogold):

Really?

OpenStudy (anonymous):

yeah, this is from connections academy right?

OpenStudy (otonogold):

Yes, Algebra 2. And I changed it to y instead of x lol

OpenStudy (anonymous):

Okay good, haha. Yeah I'm doing my algebra 2 test now, and I really hope somebody answers this because I don't want to have to work all of this out!

OpenStudy (otonogold):

Me too. :P I just want to know how to do it lol

OpenStudy (otonogold):

@AllTheLonelyKillers ?

OpenStudy (otonogold):

@hero ?

OpenStudy (anonymous):

i have no clue sorry :P

OpenStudy (otonogold):

Awl, thanks anyways (:

OpenStudy (anonymous):

welcome :3

OpenStudy (otonogold):

@peyxo I have the answers for 1-4 so far :D

OpenStudy (anonymous):

Yeah I have 1-4 too! For #5 I chose the last answer choice. Graph all the lines on a coordinate plane. For x>1, shade to the right because x is greater, for x<7, shade to the left because x is less, for y>2, shade above because y is greater, and for y<-1/3x + 6 shade below because y is less. The answer choice that best shows the area where all the shading overlaps is the last one! I'm pretty sure it's right.

OpenStudy (otonogold):

Yea, your right, I put it into wolframalpha and got D :)

OpenStudy (anonymous):

Okay awesome!

OpenStudy (otonogold):

My new math buddy ! xD

OpenStudy (anonymous):

Haha yeah I'm honestly not the best at math, but if you have any questions you can definitely ask me!

OpenStudy (otonogold):

Okay and if I find any answers , Ill make sure to let you know :)

OpenStudy (anonymous):

@OtonoGold did you finish the test? can you send me the answers?

OpenStudy (anonymous):

1. (1,2) 2. 4 3. C 4.B. x+y=700 2x+4.5y=1200 5. D 6. -4 7. 6 lbs gummy candy, 2lbs. jelly beans, 4 lbs hard candy 8. Infinitely many solutions 9. B. 4x+7y=-12 -14x+4y=4 10. D (-1,-10) 11. No solutions. 12. (1.5, 1.5) 13. Didn't do this problem. 14. (4,-1, -3)